# Kolmogorov-size particles in homogeneous and isotropic turbulence

### v1

@@ 3.3 Scale-by-scale energy budget
We now detail the influence of the particles on the organisation of the fluctuations, by studying the scale-by-scale energy budget equation. For the present case with three homogeneous directions, the energy balance reads
@@
Figure 5: Scale-by-scale energy budget for (top) $\Phi_{V}=10^{-5}$ and (bottom) $\Phi_{V}=10^{-3}$. Plots in the left column are for $\rho_{p}/\rho_{f}=5$, and those in the right for $\rho/\rho_{f}=100$. The symbols in the bottom right panel are from the simulation carried out with the coarser grid; circles are for $\Pi/\epsilon$, triangles for $\Pi_{fs}/\epsilon$ and squares for $D_{v}/\epsilon$.
Figure 5 plots the $\Pi,\Pi_{fs}$ and $D_{v}$ terms of equation 5 as a function of $\kappa$ for the four particle-laden cases investigated. For validation purposes, we also plot with symbols the terms obtained with the coarser grid for $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$. In agreement with the multiscaling behaviour shown in the energy spectra (see figure 2), the energy budgets exhibit two distinct behaviours. Energy is injected at the largest scales at a rate that equals the dissipation rate $P(0)=\epsilon$ (not shown). In the inertial range of scales $\kappa_{L}<\kappa<\kappa_{p}$, the fluid-solid coupling term is subdominant and
@@
Here the fluctuations that are produced by the fluid-solid interaction are directly dissipated by viscosity. In agreement with the energy spectrum, the range of scales where this regime holds widens as $\Phi_{V}$ and/or $\rho_{p}/\rho_{f}$ are increased.
@@ 3.4 The local structure of the flow
As shown in §3.1, Kolmogorov-size particles mainly modify the organisation of the velocity fluctuations at the smallest scales. Particles indeed modulate the energy spectrum and the structure functions for $\kappa\gtrapprox\kappa_{p}$ only. Here we characterise the smallest scales of the flow to provide new insights of the influence of the dispersed phase on the structure of the velocity fluctuations. For the sake of brevity, in this section we consider the $\Phi_{V}=10^{-3}$ cases only, and we investigate how particles modify the velocity gradient field $A_{ij}=\partial u_{j}/\partial x_{i}$. In the neighbourhood of a given point $(\bm{x}_{0},t)$, the velocity field can be approximated as $u_{i}(\bm{x},t)=u_{i}(\bm{x}_{0},t)+A_{ij}(\bm{x}_{0},t)(x_{j}-x_{0,j})+\mathcal{O}(|\bm{x}-\bm{x}_{0}|^{2})$. This linear expansion is valid in the region around $\bm{x}_{0}$, where the fluid is sufficiently smooth and the variations of $A_{ij}$ are small (Meneveau, 2011); for a turbulent flow the extent of this region is of the order of the Kolmogorov scale $\eta$. Based on these arguments, we study the influence of the particles on the smallest scales of the flow, by inspecting their effect on the $A_{ij}$ tensor.
@@
Figure 7 shows the influence of the solid phase on (top) the square of the vorticity magnitude $\omega^{2}$, i.e. the enstrophy, and on (bottom) the alignment between the vorticity $\bm{\omega}$ and the principal axes of strain. The distribution of $\omega^{2}$ shows that the tail becomes longer and the probability of large events increases due to the presence of the particles. This agrees with the enhanced flow intermittency discussed in §3.2. The tail of the distribution is longer for the $\rho_{p}/\rho_{f}=100$ case, as the velocity gradients at the particles surface are more intense because of the larger relative velocity between the particles and the surrounding fluid phase. However, even the lighter particles with $\rho_{p}/\rho_{f}=5$ are able to produce a non-negligible change of the distribution. Instead, the bottom panel of figure 7 shows that the alignment between $\bm{\omega}$ and the principal axes of strain are only slightly influenced by the presence of the particles. The results for the single-phase case perfectly overlaps with that of $\rho_{p}/\rho_{f}=5$. The presence of the heavy particles, instead, slightly reduces the alignment between $\bm{\omega}$ and $\hat{\bm{e}}_{\beta}$, as well as the anti-alignment between $\bm{\omega}$ and $\hat{\bm{e}}_{\gamma}$. Our results suggest that for $D_{p}/\eta\approx 1$ the perturbation field induced by the heavy particles is characterised by events where vorticity is more aligned with the direction of extension ($\hat{\bm{e}}_{\alpha}$) and compression ($\hat{\bm{e}}_{\gamma}$), and more anti-aligned with the intermediate eigenvector ($\hat{\bm{e}}_{\beta}$). Interestingly, this differs from what observed by Cannon et al. (2024), who found the opposite trend for particles with size in the inertial range. We presume that the difference is due to the different particle Reynolds number, which results into a different kind of wake.
@@
Figure 8: $Q$-$R$ map (top) and distributions of $Q$ (bottom left) and $R$ (bottom right) for $\Phi_{V}=10^{-3}$. For the sake of clarity the top panel show only the single-phase and the $\rho_{p}/\rho_{f}=100$ cases.
@@
Figure 9: Volumetric rendering of the (left) $Q$ and (right) $R$ fields for $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$. Orange and magenta regions are associated with large, positive values of $Q$ and $R$, while the indigo and green ones indicate large, negative values.
@@
Figure 8, plots the joint distribution of $Q$ and $R$ for $\Phi_{V}=10^{-3}$; the black solid lines denote the left- and right-Vieillefosse tails with $\Delta=0$. For completeness the bottom panels report the distributions of $Q$ (left) and $R$ (right). In absence of the particles, the $Q$-$R$ joint distribution takes a tear-drop pattern, with a clear point at the right-Vieillefosse tail with $R>0$ and $Q<0$ (Meneveau, 2011). The distribution is skewed towards positive $Q$, but rather evenly distributed among positive and negative values of $R$ (see the bottom panels). The largest probability is observed in the second and fourth quadrants, i.e. $Q<0$ $R>0$ and $Q>0$ and $R<0$. Thus, in a purely Newtonian turbulent flow there is a strong negative correlation between $Q$ and $R$, and the two most common states are vortex stretching $\omega_{i}\omega_{j}S_{ij}>0$, and biaxial strain $\alpha\beta\gamma<0$ (Betchov, 1956; Davidson, 2004). Also, the points in the $Q$-$R$ map are distributed around the origin, since the mean values of $Q$ and $R$ are zero in homogeneous flow (Nomura & Post, 1998). The presence of the particles enlarges the range of possible $Q$ and $R$ values, in agreement with the increase of the probability of events with intense velocity gradients associated with the boundary conditions at the particles’ surface. In particular, Kolmogorov-size particles mainly favour events that lay in the first and third quadrant, resulting into a joint distribution that is more symmetric with respect to an inversion of the $R$ axis (see the bottom right panel in figure 8). Compared to the unladen case, particles mainly promote events with axial strain $\alpha\beta\gamma>0$ ($R<0$ and large negative $Q<0$) and with vortex compression $\omega_{i}\omega_{j}S_{ij}<0$ ($R>0$ and $Q>0$). The probability of events with $R<0$ and $Q<0$ is particularly enhanced, as visualised in the top panel of figure 8 by the occurrence of a point at the left-Vieillefosse tail. This is consistent with the increase of the probability of events with $s^{*}=-1$ shown in figure 6. A visual investigation of the $Q$ and $R$ fields around the particles (figure 9) helps to explain this effect by highlighting the local contribution of the particles. By comparing the two fields in the surroundings of the particles, we observe that they are both almost axisymmetric along the axis aligned with their travelling direction. However, across the velocity-normal median plane, $Q$ is symmetric, while $R$ is antisymmetric. This relation between the two fields implies contributions across all four quadrants of the $Q$-$R$ distribution. The effect, however, is particularly apparent in the third quadrant, as the region is otherwise not explored by the single-phase flow. In particular, regions of $Q<0$ and $R<0$ are associated with axial strain and appear to be found at one of the two ends of the particles’ along their travelling direction. Consistently with the above-mentioned symmetries, at the opposite end of the particles, a region of biaxial strain ($Q<0$ and $R>0$) is found.
@@
This scenario only partially agrees with the results of Schneiders et al. (2017) for decaying homogeneous isotropic turbulence laden with Kolmogorov-size particles. They also found that particles favour events with axial strain ($Q<0$ and $R<0$) and biaxial strain ($Q<0$ and $R>0$), as shown by the occurrence of a point at the left-Vieillefosse tail and by the more pronounced right-Vieillefosse tail in figure 20 of their paper. However, they did not report an increase of the probability of events with $Q>0$ like in the present case. It is worth mentioning, however, that they do report a global increase of the vortex stretching. Compared to larger particles with size in the inertial range of scales, the scenario is completely different: in fact, Cannon et al. (2024) found that when large particles are added both $Q$ and $R$ are reduced. Besides energising the small scales, large particles indeed behave also as obstacles for the large flow structures, largely weakening thus the energy content at the large scales. Cannon et al. (2024) only observed an increase of the probability in the strain-dominated region (that we also observe), which is an indication of the intense dissipation regions that arise around the particles.
Figure 10: $-Q_{S}$-$Q_{W}$ map (top) and distributions of $Q_{S}$ (bottom left) and $Q_{W}$ (bottom right) for $\Phi_{V}=10^{-3}$. For the sake of clarity the top panel show only the single-phase and the $\rho_{p}/\rho_{f}=100$ cases.
@@
These invariants are related with the fluid dissipation $\epsilon=-4\nu Q_{S}$ and with the fluid enstrophy $\omega^{2}=4Q_{W}$. Therefore, the $Q_{S}$-$Q_{W}$ joint distribution determines whether the flow is dominated by dissipation (extensional dominated regions with $Q_{S}>Q_{W}$) or by enstrophy (rigid rotation regions with $Q_{W}>Q_{S}$). In shear dominated regions, dissipation and enstrophy balance and $-Q_{S}=Q_{W}$ (Soria et al., 1994). For simplicity, we follow Truesdell (1954) and introduce $\mathcal{K}=(-Q_{W}/Q_{S})^{1/2}$; when $\mathcal{K}=0$ the flow is extension dominated, when $\mathcal{K}=\infty$ the flow is dominated by rigid rotation events, and when $\mathcal{K}=1$ the rotation and the stretching are equal, as typical of vortex sheets and shear layers. In the unladen case, events with $Q_{W}>-Q_{S}$ ($\mathcal{K}\rightarrow\infty$) are more frequent, meaning that for purely Newtonian turbulence the flow is mainly dominated by rigid rotations; see also the distributions of $Q_{S}$ and $Q_{W}$ in the bottom panels of figure 10. In the presence of the particles the scenario slightly changes. A first observation is that the probability of events with large $Q_{S}$ (or $\mathcal{K}=0$) increases, indicating that the perturbation field induced by these small particles is extensional dominated. A second observation is that the solid phase favours also events with $-Q_{S}\approx Q_{W}$ ($\mathcal{K}\approx 1$), that agrees with the presence of the shear layers induced by the presence of the particles.
@@ 4.1 Lagrangian velocity increments
Figure 11: Probability density function of the Lagrangian velocity increments of the particles $\delta_{\tau}u_{p}=u_{p}(t+\tau)-u_{p}(t)$ for (left) $\tau=0.2\tau_{\eta}$, and (right) $\tau=2\tau_{\eta}$. The data are for the case with $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=5$ and $\rho_{p}/\rho_{f}=100$
Figure 12: Probability density function of the Lagrangian velocity increments of the particles $\delta_{\tau}u_{p}=u_{p}(t+\tau)-u_{p}(t)$ for (top) $\tau=0.2\tau_{\eta}$, (centre) $\tau=2\tau_{\eta}$, and (bottom) $\tau=30\tau_{\eta}$. Left panels are for $\rho_{p}/\rho_{f}=5$, while the right ones for $\rho_{p}/\rho_{f}=100$ The black solid lines are the results of the PP-DNS simulations.
The Lagrangian statistics of the particles motion are of fundamental importance in the understanding of transport and mixing. In order to investigate this, we study the Lagrangian velocity increments, defined here as $\delta_{\tau}u_{p,i}=u_{p,i}(t+\tau)-u_{p,i}(t)$, with $u_{p,i}(t)$ being the instantaneous velocity of a particle along direction $i$ at time $t$. The symmetries of the present problem make the statistics of $\delta_{\tau}u_{p,i}$ independent on both $t$ and $i$; for simplicity hereafter we drop the $i$ index. Figures 11 and 12 describe the particle dynamics at different time scales, by plotting $\delta_{\tau}u_{p}$ for different values of the time scale $\tau$ in the $0.2\tau_{\eta}\leq\tau\leq 30\tau_{\eta}$ range, where $\tau_{\eta}=(\nu/\epsilon)^{1/2}$ is the Kolmogorov time scale. For small time scales, the velocity increment provides information about the particle acceleration, i.e. $\delta_{\tau}u_{p}\sim a_{p}\tau$. A first observation is that the distributions are symmetric, in agreement with the symmetries of the flow. The probability density function of $\delta_{\tau}u_{p}$ continuously deforms from the Gaussian at large time scales (see $\tau\approx 30\tau_{\eta}$) to the development of stretched exponential tails at dissipative time scales (see $\tau\approx 0.2\tau_{\eta}$), which are the statistical signature of an intermittent Lagrangian dynamics; see Mordant et al. (2001), La Porta et al. (2001) and Chevillard et al. (2003) for small tracers and Qureshi et al. (2007) for finite-size neutrally buoyant particles. The wide stretched exponential tails for the smallest $\tau$ show that the finite-size particles with $D_{p}\approx\eta$ experience very high acceleration events, with a probability which is higher than Gaussian, similarly to what was found for small tracers by La Porta et al. (2001) and for finite-size particles by Qureshi et al. (2007).
@@
Figure 13: Evolution of the excess kurtosis factor $\mathcal{K}_{\delta_{\tau}u_{p}}(\tau)=\!\left\langle{\delta_{\tau}u_{p}^{4}}\right\rangle\!/\!\left\langle{\delta_{\tau}u_{p}^{2}}\right\rangle\!^{2}-3$ for the distributions of the time increments of the particle velocity. Left: $\rho_{p}/\rho_{f}=5$. Right: $\rho_{p}/\rho_{f}=100$. The empty circles in the right panel are from the simulation with the coarser grid at $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$, and are shown for validation purpose.
The left panels of figure 12 show that for $\rho_{p}/\rho_{f}=5$ the distributions obtained by means of PP-DNS and PR-DNS almost perfectly overlap for all time scales $\tau$: for light particles the complete MRG equation properly predicts the Lagrangian intermittency of the particles dynamics. For heavier particles the match between the PP-DNS and the PR-DNS is rather good at large time scales, but differences are observed at small $\tau$, where the PP-DNS does not predict the large tails for $\tau\lessapprox 2\tau_{\eta}$. As discussed above, these extreme events are associated with the flow modulation which is not modelled in our PP-DNS. The comparison between the PP-DNS and PR-DNS results is further detailed in figure 13 where the evolution of the $\delta_{\tau}u_{p}$ distribution with $\tau$ is quantified by means of the excess kurtosis $\mathcal{K}_{\delta_{\tau}u_{p}}(\tau)=\!\left\langle{\delta_{\tau}u_{p}^{4}}\right\rangle\!/\!\left\langle{\delta_{\tau}u_{p}^{2}}\right\rangle\!^{2}-3$. At large scales $\mathcal{K}_{\delta_{\tau}u_{p}}\approx 0$ in agreement with the Gaussian-like shape of the distribution, while it steeply increases at small scales. For $\rho_{p}/\rho_{f}=5$ the good agreement between the PP-DNS and the PR-DNS is again clear, with a small deviation for the $\Phi_{V}=10^{-3}$ case, which is due to the non-zero flow modulation. For $\rho_{p}/\rho_{f}=100$ the agreement is good at large scales, while the three curves substantially deviate for small $\tau$, in agreement with the larger tails of $\delta_{\tau}u_{p}$ found in the PR-DNS.

### v2

@@ 3.3 Scale-by-scale energy budget
We now detail the influence of the particles on the organisation of the fluctuations, by studying the scale-by-scale energy budget equation and characterising thus the dominant energetic mechanisms in the different regimes identified with the energy spectra (see figure 2). For the present case with three homogeneous directions, the energy balance reads
@@
Figure 5: Scale-by-scale energy budget for (top) $\Phi_{V}=10^{-5}$ and (bottom) $\Phi_{V}=10^{-3}$. Plots in the left column are for $\rho_{p}/\rho_{f}=5$, and those in the right for $\rho/\rho_{f}=100$. The production term $P$ acts at the largest scales $\kappa/\kappa_{L}\leq 1$ only, being $P=0$ for $\kappa/\kappa_{L}\geq 2$ (not visible as the $y$ axes is in log scale). The filled symbols in the bottom right panel are from the simulation carried out with the coarser grid; circles are for $\Pi/\epsilon$, triangles for $\Pi_{fs}/\epsilon$ and squares for $D_{v}/\epsilon$.
Figure 5 plots the terms of equation 5 as a function of $\kappa$ for the four particle-laden cases investigated. For validation purposes, we also plot with filled symbols the terms obtained with the coarser grid for $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$. In agreement with the multiscaling behaviour shown in the energy spectra (see figure 2), the energy budgets exhibit two distinct behaviours. Energy is injected at the largest scales at a rate that equals the dissipation rate $P(0)=\epsilon$. In the inertial range of scales $\kappa_{L}<\kappa<\kappa_{p}$, the fluid-solid coupling term is subdominant and
@@
Here the fluctuations that are produced by the fluid-solid interaction are (on average) directly dissipated by viscosity and are not transferred among scales. In agreement with the energy spectrum, the range of scales where this regime holds widens as $\Phi_{V}$ and/or $\rho_{p}/\rho_{f}$ are increased.
@@ 3.4 The local structure of the flow
As shown in §3.1, Kolmogorov-size particles mainly modify the organisation of the velocity fluctuations at the smallest scales. Particles indeed modulate the energy spectrum and the structure functions for $\kappa\gtrapprox\kappa_{p}$ only. Here we characterise the smallest scales of the flow to provide new insights of the influence of the dispersed phase on the structure of the velocity fluctuations. For the sake of brevity, in this section we consider the $\Phi_{V}=10^{-3}$ cases only, and we investigate how particles modify the velocity gradient field $A_{ij}=\partial u_{j}/\partial x_{i}$. In the neighbourhood of a given point $(\bm{x}_{0},t)$, indeed, the velocity field can be approximated as $u_{i}(\bm{x},t)=u_{i}(\bm{x}_{0},t)+A_{ij}(\bm{x}_{0},t)(x_{j}-x_{0,j})+\mathcal{O}(|\bm{x}-\bm{x}_{0}|^{2})$. This linear expansion is valid in the region around $\bm{x}_{0}$, where the fluid is sufficiently smooth and the variations of $A_{ij}$ are small (Meneveau, 2011); for a turbulent flow the extent of this region is of the order of the Kolmogorov scale $\eta$. Based on these arguments, we study the influence of the particles on the smallest scales of the flow, by inspecting their effect on the $A_{ij}$ tensor.
@@
Figure 7 shows the influence of the solid phase on (top) the square of the vorticity magnitude $\omega^{2}$, i.e. the enstrophy, and on (bottom) the alignment between the vorticity $\bm{\omega}$ and the principal axes of strain. The distribution of $\omega^{2}$ shows that the tail becomes longer and the probability of large events increases due to the presence of the particles. This agrees with the enhanced flow intermittency discussed in §3.2. The tail of the distribution is longer for the $\rho_{p}/\rho_{f}=100$ case, as the velocity gradients at the particles surface are more intense because of the larger relative velocity between the particles and the surrounding fluid phase. However, even the lighter particles with $\rho_{p}/\rho_{f}=5$ are able to produce a non-negligible change of the distribution. Instead, the bottom panel of figure 7 shows that the alignment between $\bm{\omega}$ and the principal axes of strain are only slightly influenced by the presence of the particles. The results for the single-phase case perfectly overlaps with that of $\rho_{p}/\rho_{f}=5$. The presence of the heavy particles, instead, slightly reduces the alignment between $\bm{\omega}$ and $\hat{\bm{e}}_{\beta}$, as well as the anti-alignment between $\bm{\omega}$ and $\hat{\bm{e}}_{\gamma}$. Our results suggest that for $D_{p}/\eta\approx 1$ the perturbation field induced by the heavy particles is characterised by events where vorticity is more aligned with the direction of extension ($\hat{\bm{e}}_{\alpha}$) and compression ($\hat{\bm{e}}_{\gamma}$), and more anti-aligned with the intermediate eigenvector ($\hat{\bm{e}}_{\beta}$). Interestingly, this differs from what observed by Cannon et al. (2024), who found the opposite trend for particles with size in the inertial range. We presume that the difference is due to the different particle Reynolds number, which results in a different kind of deformation of the local fluid flow.
@@
Figure 8: Top: $Q$-$R$ map for the single-phase (left) and $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$ (right). For both panels the $11$ isolines have a logarithmic spacing between $5\times 10^{-5}$ and $10^{2}$. Bottom: distribution of $Q$ (left) and $R$ (right) for $\Phi_{V}=10^{-3}$.
@@
Figure 9: Volumetric rendering of the (left) $Q$ and (right) $R$ fields for $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$. Orange and magenta regions are associated with large, positive values of $Q$ and $R$, i.e. $0.2072\lessapprox Q\tau_{\eta}^{2}\lessapprox 2.072$ and $0.2109\lessapprox R\tau_{\eta}^{3}\lessapprox 0.4218$. Indigo and green regions indicate large, negative values of $Q$ and $R$, i.e. $-2.072\lessapprox Q\tau_{\eta}^{2}\lessapprox-0.2072$ and $-0.4218\lessapprox R\tau_{\eta}^{3}\lessapprox-0.2109$. The particle travelling direction relative to the local fluid velocity in a shell of radius $R_{sh}=5$ is indicated for each particle by a pointer. See figure 10 for a schematic representation of the flow.
Figure 10: Schematic representation of the distributions of $Q$ and $R$ around the particles based on figure 9. The flow, represented by continuous stream lines, is incoming from the left in the particle’s reference frame.
@@
Figure 8, plots the joint distribution of $Q$ and $R$ for the $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$ case (top right) and for the single phase (top left); the black, thin, solid (curved) lines denote the left- and right-Vieillefosse tails with $\Delta=0$. For completeness the bottom panels report the distributions of $Q$ (left) and $R$ (right). In absence of the particles, the $Q$-$R$ joint distribution takes a tear-drop pattern, with a clear point at the right-Vieillefosse tail with $R>0$ and $Q<0$ (Meneveau, 2011). The distribution is skewed towards positive $Q$, but rather evenly distributed among positive and negative values of $R$ (see the bottom panels). The largest probability is observed in the second and fourth quadrants, i.e. $Q<0$ and $R>0$, and $Q>0$ and $R<0$. Thus, in a purely Newtonian turbulent flow there is a strong negative correlation between $Q$ and $R$, and the two most common states are vortex stretching $\omega_{i}\omega_{j}S_{ij}>0$, and biaxial strain $\alpha\beta\gamma<0$ (Betchov, 1956; Davidson, 2004). Also, the points in the $Q$-$R$ map are distributed around the origin, since the mean values of $Q$ and $R$ are zero in homogeneous flow (Nomura & Post, 1998). The presence of the particles enlarges the range of possible $Q$ and $R$ values, in agreement with the increase of the probability of events with intense velocity gradients associated with the boundary conditions at the particles’ surface. In particular, Kolmogorov-size particles mainly favour events that lay in the first and third quadrant, resulting in a joint distribution that is more symmetric with respect to an inversion of the $R$ axis (see the bottom right panel in figure 8). Compared to the unladen case, particles mainly promote events with axial strain $\alpha\beta\gamma>0$ ($R<0$ and large negative $Q<0$) and with vortex compression $\omega_{i}\omega_{j}S_{ij}<0$ ($R>0$ and $Q>0$). The probability of events with $R<0$ and $Q<0$ is particularly enhanced, as visualised in the top right panel of figure 8 by the occurrence of a point at the left-Vieillefosse tail. This is consistent with the increase of the probability of events with $s^{*}=-1$ shown in figure 6. A visual investigation of the $Q$ and $R$ fields around the particles (figure 9) helps to explain this effect by highlighting the local contribution of the particles. By comparing the two fields in the surroundings of the particles, we observe that they are both almost axisymmetric along the axis aligned with their travelling direction. However, across the velocity-normal median plane, $Q$ is symmetric, while $R$ is antisymmetric. This relation between the two fields implies contributions across all four quadrants of the $Q$-$R$ distribution (figure 8). The effect, however, is particularly apparent in the third quadrant, as the region is otherwise not explored by the single-phase flow. In particular, regions of $Q<0$ and $R<0$ are associated with axial strain and appear to be found at the downstream end of the particles’ along their travelling direction. Consistently with the above-mentioned symmetries, a region of biaxial strain ($Q<0$ and $R>0$) is found at the upstream end of the particles. This is conveniently visualised in the schematic of figure 10.
@@
This scenario only partially agrees with the results of Schneiders et al. (2017) for decaying homogeneous isotropic turbulence laden with Kolmogorov-size particles. They also found that particles favour events with axial strain ($Q<0$ and $R<0$) and biaxial strain ($Q<0$ and $R>0$), as shown by the occurrence of a point at the left-Vieillefosse tail and by the more pronounced right-Vieillefosse tail in figure 20 of their paper. However, they did not report an increase of the probability of events with $Q>0$ like in the present case. It is worth mentioning, however, that they do report a global increase of the vortex stretching. A possible explanation of this difference is the lower Reynolds number they considered ($Re_{\lambda}\approx 79$ at the initial time) that does not ensure a proper separation of scales, in particular at larger times when turbulence decays. Compared to larger particles with size in the inertial range of scales, the scenario is completely different: in fact, Cannon et al. (2024) found that when large particles are added both $Q$ and $R$ are reduced. Besides energising the small scales, large particles indeed behave also as obstacles for the large flow structures, largely weakening thus the energy content at the large scales. Cannon et al. (2024) only observed an increase of the probability in the strain-dominated region (that we also observe), which is an indication of the intense dissipation regions that arise around the particles.
Figure 11: $-Q_{S}$-$Q_{W}$ map for the single-phase case (left) and $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$ (right). For both panels the $20$ isolines have a logarithmic spacing between $10^{-7}$ and $10^{1}$.
@@
These invariants are related with the fluid dissipation $\epsilon=-4\nu Q_{S}$ and with the fluid enstrophy $\omega^{2}=4Q_{W}$. Therefore, the $Q_{S}$-$Q_{W}$ joint distribution determines whether the flow is dominated by dissipation (extensional dominated regions with $Q_{S}>Q_{W}$) or by enstrophy (rigid rotation regions with $Q_{W}>Q_{S}$). In shear dominated regions, dissipation and enstrophy balance and $-Q_{S}=Q_{W}$ (Soria et al., 1994). For simplicity, we follow Truesdell (1954) and introduce $\mathcal{K}=(-Q_{W}/Q_{S})^{1/2}$; when $\mathcal{K}=0$ the flow is extension dominated, when $\mathcal{K}=\infty$ the flow is dominated by rigid rotation events, and when $\mathcal{K}=1$ the rotation and the stretching are equal, as typical of vortex sheets and shear layers. In the unladen case, events with $Q_{W}>-Q_{S}$ ($\mathcal{K}\rightarrow\infty$) are more frequent, meaning that for purely Newtonian turbulence the flow is mainly dominated by rigid rotations. In the presence of the particles the scenario slightly changes. A first observation is that the probability of events with large $Q_{S}$ (or $\mathcal{K}=0$) increases, indicating that the perturbation field induced by these small particles is extensional dominated. A second observation is that the solid phase favours also events with $-Q_{S}\approx Q_{W}$ ($\mathcal{K}\approx 1$), that agrees with the presence of the shear layers induced by the presence of the particles.
@@ 4.1 Lagrangian velocity increments
Figure 12: Probability density function of the Lagrangian velocity increments of the particles $\delta_{\tau}u_{p}$ for (left) $\tau=0.2\tau_{\eta}$, and (right) $\tau=2\tau_{\eta}$. The data are for the case with $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=5$ and $\rho_{p}/\rho_{f}=100$
[…]
Figure 13: Probability density function of the Lagrangian velocity increments of the particles $\delta_{\tau}u_{p}$for (top) $\tau=0.2\tau_{\eta}$, (centre) $\tau=2\tau_{\eta}$, and (bottom) $\tau=30\tau_{\eta}$. Left panels are for $\rho_{p}/\rho_{f}=5$, while the right ones for $\rho_{p}/\rho_{f}=100$ The black solid lines are the results of the PP-DNS simulations.
The Lagrangian statistics of the particles motion are of fundamental importance in the understanding of transport and mixing. In order to investigate this, we study the Lagrangian velocity increments, defined here as $\delta_{\tau}u_{p,i}=u_{p,i}(t+\tau)-u_{p,i}(t)$, with $u_{p,i}(t)$ being the instantaneous velocity of a particle along direction $i$ at time $t$. The symmetries of the present problem make the statistics of $\delta_{\tau}u_{p,i}$ independent of both $t$ and $i$; for simplicity hereafter we drop the $i$ index. Figures 12 and 13 describe the particle dynamics at different time scales, by plotting $\delta_{\tau}u_{p}$ for different values of the time scale $\tau$ in the $0.2\tau_{\eta}\leq\tau\leq 30\tau_{\eta}$ range, where $\tau_{\eta}=(\nu/\epsilon)^{1/2}$ is the Kolmogorov time scale. For small time scales, the velocity increment provides information about the particle acceleration, i.e. $\delta_{\tau}u_{p}\sim a_{p}\tau$. A first observation is that the distributions are symmetric, in agreement with the symmetries of the flow. The probability density function of $\delta_{\tau}u_{p}$ continuously deforms from the Gaussian at large time scales (see $\tau\approx 30\tau_{\eta}$) to the development of stretched exponential tails at dissipative time scales (see $\tau\approx 0.2\tau_{\eta}$), which are the statistical signature of an intermittent Lagrangian dynamics; see Mordant et al. (2001), La Porta et al. (2001) and Chevillard et al. (2003) for small tracers and Qureshi et al. (2007) for finite-size neutrally buoyant particles. The wide stretched exponential tails for the smallest $\tau$ show that the finite-size particles with $D_{p}\approx\eta$ experience very high acceleration events, with a probability which is higher than Gaussian, similarly to what was found for small tracers by La Porta et al. (2001) and for finite-size particles by Qureshi et al. (2007).
@@
Figure 14: Evolution of the excess kurtosis factor $\mathcal{K}_{\delta_{\tau}u_{p}}(\tau)$ for the distributions of the time increments of the particle velocity. Left: $\rho_{p}/\rho_{f}=5$. Right: $\rho_{p}/\rho_{f}=100$. The empty circles in the right panel are from the simulation with the coarser grid at $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$, and are shown for validation purpose.
The left panels of figure 13 show that for $\rho_{p}/\rho_{f}=5$ the distributions obtained by means of PP-DNS and PR-DNS almost perfectly overlap for all time scales $\tau$: for light particles the complete MRG equation properly predicts the Lagrangian intermittency of the particles dynamics. For heavier particles the match between the PP-DNS and the PR-DNS is rather good at large time scales, but differences are observed at small $\tau$, where the PP-DNS does not predict the large tails for $\tau\lessapprox 2\tau_{\eta}$. As discussed above, these extreme events are associated with the flow modulation which is not modelled in our PP-DNS. The comparison between the PP-DNS and PR-DNS results is further detailed in figure 14 where the evolution of the $\delta_{\tau}u_{p}$ distribution with $\tau$ is quantified by means of the excess kurtosis $\mathcal{K}_{\delta_{\tau}u_{p}}(\tau)=\!\left\langle{\delta_{\tau}u_{p}^{4}}\right\rangle\!/\!\left\langle{\delta_{\tau}u_{p}^{2}}\right\rangle\!^{2}-3$. At large scales $\mathcal{K}_{\delta_{\tau}u_{p}}\approx 0$ in agreement with the Gaussian-like shape of the distribution, while it steeply increases at small scales. For $\rho_{p}/\rho_{f}=5$ the good agreement between the PP-DNS and the PR-DNS is again clear, with a small deviation for the $\Phi_{V}=10^{-3}$ case, which is due to the non-zero flow modulation. For $\rho_{p}/\rho_{f}=100$, instead, the agreement is good at large scales, while the three curves substantially deviate for small $\tau$, accordingly with the larger tails of $\delta_{\tau}u_{p}$ found in the PR-DNS.